# Comparing human and LLM politeness strategies in free production

### v1

@@ Abstract
Polite speech poses a fundamental alignment challenge for large language models (LLMs). Humans deploy a rich repertoire of linguistic strategies to balance informational and social goals – from positive approaches that build rapport (compliments, expressions of interest) to negative strategies that minimize imposition (hedging, indirectness). We investigate whether LLMs employ a similarly context-sensitive repertoire by comparing human and LLM responses in both constrained and open-ended production tasks. We find that larger models ($\geq$70B parameters) successfully replicate key preferences from the computational pragmatics literature, and human evaluators surprisingly prefer LLM-generated responses in open-ended contexts. However, further linguistic analyses reveal that models disproportionately rely on negative politeness strategies even in positive contexts, potentially leading to misinterpretations. While modern LLMs demonstrate an impressive handle on politeness strategies, these subtle differences raise important questions about pragmatic alignment in AI systems.
@@ 1 Introduction
Speakers do not always say exactly what they mean. For example, we might say a friend’s poem “wasn’t terrible” rather than saying “it was bad” to avoid hurting their feelings Yoon et al. (2020), or just compliment specific elements that we liked without mentioning other elements Brown and Levinson (1987); Goffman (1967); Pinker et al. (2008). These kinds of politeness strategies allow speakers to balance competing goals, conveying accurate information while maintaining positive relationships Hill et al. (1986); Leech (2014). As large language models (LLMs) are increasingly deployed in open-ended interactions across sensitive social domains like healthcare and education, their ability to appropriately use and understand polite language remains an important alignment challenge.
Politeness theory provides a valuable framework for addressing these questions. Seminal work by Brown and Levinson (1987) distinguishes between positive politeness strategies that affirm the listener (compliments, expressions of interest) and negative politeness strategies that minimize imposition (hedging, indirectness). This distinction is crucial because different contexts call for different strategies, and mismatches can lead to communication breakdowns. If an AI system employs distancing, hedge-filled language (“I’m somewhat concerned that this approach might not be optimal”) in contexts where human speakers would use rapport-building strategies (“I love your creativity here, and wonder if we could build on it by…”), users may perceive the system as cold, insincere, or lacking genuine engagement–even when its literal content is appropriate.
This pragmatic misalignment represents a critical gap in our understanding of LLMs as social agents. While considerable attention has been paid to whether models can recognize politeness or generate polite language in constrained settings, effective social interaction depends not just on understanding politeness norms in the abstract but on actively selecting and applying appropriate strategies from a diverse linguistic repertoire. Human speakers navigate this complexity intuitively, deploying hedging, elaboration, indirect speech acts, and numerous other strategies to balance competing communicative goals in context-sensitive ways. To fully understand LLMs’ grasp of politeness strategies, we need to examine whether they exhibit similar patterns of strategy selection and deployment across different contexts. This requires moving beyond limited-choice evaluations to examine open-ended language generation, where models have access to the full range of linguistic choices.
In this work, we investigate polite speech generation in both humans and LLMs, making the following contributions:
@@
Our results reveal that while LLMs have acquired important aspects of human-like pragmatic competence in polite language production – enough to be preferred by human evaluators – they also show systematic differences in strategy deployment that raise intriguing questions about the mechanisms underlying their social language capabilities. In particular, we find that models disproportionately rely on negative politeness strategies (minimizing imposition) even in contexts where humans prefer positive politeness strategies (building rapport), suggesting important differences in how these systems navigate social interactions.  [^1] Data & Code: https://github.com/haoranzhao419/politeness-speech-production
@@ 2.3 Polite Language Generation in LLMs
Work on generating polite language in AI systems represents a smaller but growing research area. Early approaches focused on style transfer, with systems like those developed by Niu and Bansal (2018) demonstrating that neural models could transform neutral text into more polite versions through specific syntactic transformations. Subsequent work explored paraphrasing to increase politeness Fu et al. (2020), politeness-focused style transfer Madaan et al. (2020), and creating polite chatbots Mukherjee et al. (2023). However, these systems typically focused on surface-level transformations rather than strategic deployment of politeness based on contextual factors. As noted in a recent survey Priya et al. (2024), existing approaches to polite language generation have predominantly emphasized isolated features (hedging expressions, please markers, specific lexical choices) rather than examining the full repertoire of politeness strategies and how they’re selected based on communicative context. This leaves a significant gap in our understanding of whether LLMs can approximate the context-sensitivity that characterizes human politeness. Our work addresses this gap by directly comparing politeness strategies in humans and LLMs across varying communicative goals, examining whether models align with human preferences for positive versus negative politeness strategies in different contexts.
@@ 3 Experiment 1: Constrained settings
Table 1: Comparison of LLM response patterns with humans (Yoon et al., 2020) and between default goal and other communicative goals (using the “multi-choice-original" prompting strategy and Spearman correlation).
To what extent are LLMs sensitive to the goals that give rise to politeness in human speech? To address this question, we first examined whether LLMs could reproduce the patterns of goal-sensitive language use reported by Yoon et al. (2020). Their study provided empirical evidence for a computational model of politeness where speakers strategically balance informational accuracy with social goals. Most notably, they found that when giving negative feedback, humans often deploy negation (e.g., “wasn’t terrible” rather than “bad”) (see also Gotzner and Scontras, 2024).
@@
We tested a range of open-source (8B-72B parameters) and closed-source models using two prompting strategies: an “original” strategy that presented scenarios verbatim, and a “persona” variant that systematically varied speaker characteristics (e.g., gender, occupation, background) to better approximate the diversity in the population of human participants Murthy et al. (2025); He-Yueya et al. (2024). For each model and prompting strategy, we sampled 30 responses with temperature $\tau=1.0$ per scenario (see Appendix A for details).
@@ 4.1 Methods
We used the scenarios from Yoon et al. (2020), preserving the same performance ratings (0-3 hearts) and communicative goals (informative, social, both, default). We collected 3 open-ended responses per scenario from 156 human participants via Prolific (each responding to 4 distinct scenarios) and three responses from LLMs that performed well in our first experiment: GPT-4o, Claude-3.5-Sonnet, and Llama-3.3-70B-Instruct. Models were instructed to “keep responses short and concise” to ensure comparable length with human responses. To assess preferences for these responses, we then conducted a two-alternative forced-choice evaluation with 156 human evaluators, each viewing four different scenarios. Evaluators made five judgments per scenario: (1) comparing human vs. LLM responses, (2-3) comparing goal-congruent vs. goal-incongruent responses for both sources, and (4-5) comparing rating-congruent vs. rating-incongruent responses for both sources. We randomized presentation order and ensured evaluators saw responses from different sources across blocks (see Appendix B for full details).
@@ 5.2 Word usage patterns
To better understand differences between human and LLM responses, we analyzed unigram distributions using Pointwise Mutual Information (PMI) and Jensen-Shannon Divergence (JSD). First, examining unigrams with highest PMI, we found that human responses more frequently incorporated casual language and expressions (e.g., “awesome," “great"), whereas LLM-generated responses tended toward more formal linguistic choices (e.g., “fabulous," “excellent"). Both groups effectively employed personalization as a politeness strategy, such as directly mentioning the performer’s name (e.g., “Your app is pretty good, Henry!"). Additionally, both humans and LLMs adapted their lexical choices based on context, with minimal overlap in high-PMI words across different goals and ratings.
@@ 5.3 Annotated politeness strategies
To obtain a comprehensive picture of the politeness strategies employed in human and LLM responses, we conducted a detailed annotation using the politeness framework from Brown and Levinson (1987), supplemented by markers from Danescu-Niculescu-Mizil et al. (2013). This framework distinguishes four broad categories of politeness strategies with many subtypes (see Appendix Table 10): positive politeness (e.g. compliments and expressions of interest), negative politeness (e.g. hedging and indirectness), off-record strategies (indirect hints that maintain plausible deniability), and bald-on-record strategies (direct statements without politeness). We used LLMs (GPT-4.1 and Claude-3.7-Sonnet) as annotators, following best practices Tan et al. (2024). Annotations were manually verified and corrected where necessary. This approach allowed us to identify specific politeness markers and their associated strategies across the dataset. Appendix C.3 provides full details on the annotation process and framework.
Strategy Distribution.
As shown in Figure 3, both humans and LLMs rely primarily on positive politeness (rapport-building) and negative politeness (minimizing imposition) strategies, with relatively low use of off-record and bald-on-record approaches. However, a key difference emerged in strategy selection patterns: while both humans and LLMs increased their use of positive politeness strategies as ratings increased, LLMs showed systematically higher use of negative politeness strategies even in positive contexts (higher ratings), where humans tended to reduce such strategies. These strategy distributions were significantly different under a permutation test (JSD = 0.023, $p<0.001$). The overreliance on negative politeness strategies in positive contexts represents a fundamental misalignment with human communication patterns. It suggests LLMs may have been trained to prioritize non-imposition over positive affirmation, potentially reflecting alignment procedures that emphasize harm reduction over more natural social interactions.
@@
6 Conclusion
@@
While LLMs demonstrate impressive pragmatic competence in politeness, we find that they also systematically differ from humans in how they deploy different strategies. Most notably, LLMs over-rely on negative politeness strategies even in positive contexts where humans prefer rapport-building approaches. This mismatch persists despite LLM responses being consistently preferred by human evaluators in open-ended settings, suggesting that surface-level preferences may not fully capture important pragmatic differences. The contrast between constrained and open-ended tasks further highlights how LLMs leverage a richer repertoire of strategies when given more expressivity, such as the “while compliment, suggestion” constructions we observed in low-rating scenarios.
The distributional differences in politeness strategy deployment may have significant implications for human-AI communication. When AI systems consistently employ distancing strategies even when delivering positive feedback, this contextual mismatch could lead to pragmatic misinterpretations, where humans interpret hedged positive feedback as more negative than intended. Such patterns may also reduce social presence by making AI communication feel unnecessarily formal and detached, ultimately diminishing information transfer if humans expend additional cognitive resources parsing unnecessarily indirect language. These consequences extend beyond mere stylistic differences to potentially impact the fundamental goals of communication: accurate information transfer and relationship maintenance.
Our results underscore the need to investigate not just whether AI systems can produce polite language, but whether their specific patterns of politeness strategies facilitate or hinder effective communication. For example, future work should directly test how differences in strategy deployment affect human listeners’ ability to recover underlying information and speakers’ intentions, particularly in extended multi-turn conversations and across diverse cultural contexts. The systematic divergence we observed raises a concern that even if LLM responses are preferred in isolation, their departure from human pragmatic patterns may lead to miscommunications in real-world interactions where precise understanding of intentions and attitudes is crucial. These findings could inform future LLM training approaches to better align with human pragmatic patterns, particularly in balancing positive and negative politeness strategies.
Finally, our results have implications for rational models of politeness in computational linguistics: given the richer space of strategies exposed in open-ended production, it will be essential to develop richer computational models that can probe the implicit tradeoffs guiding LLM behavior. For instance, we observed that LLMs commonly use constructions like “compliment but suggestion” in low-rating scenarios, a strategy that balances kindness with informativeness in ways that cannot be accomplished via simple negation. But, if we could be so bold as to make one suggestion, we’d greatly appreciate it if future work could help systems learn when to use positive strategies rather than minimize imposition.
@@
While our work gave a comprehensive picture of comparing the polite language use in humans and LLMs, there are still limitations that could be addressed in future work. First, throughout our analyses, we still cannot answer the question of what makes human evaluators prefer the responses they prefer, as all our analyses showed very minimal differences between preferred and non-preferred responses. One guess is that even ratings and goals are made very clear in the provided scenarios, human evaluators still may not pay enough attention to, and optionally omit this information, instead, they tend to pick whichever one in the given pair that sounds nicer. Future research, for example, testing LLMs as evaluators and comparing LLM-as-evaluator preference results with humans, could give us more insight into this question. Additionally, as our results show that LLMs are still not quite human-like in picking the right politeness strategies in a context-sensitive way, future research on how to develop computational methods and algorithms to make LLMs better at polite language use and as social agents will be necessary.

### v2

@@ Abstract
Polite speech poses a fundamental alignment challenge for large language models (LLMs). Humans deploy a rich repertoire of linguistic strategies to balance informational and social goals – from positive approaches that build rapport (compliments, expressions of interest) to negative strategies that minimize imposition (hedging, indirectness). We investigate whether LLMs employ a similarly context-sensitive repertoire by comparing human and LLM responses to English-language scenarios in both constrained and open-ended production tasks. We find that larger models ($\geq$70B parameters) successfully replicate key effects from the computational pragmatics literature, and human evaluators prefer LLM-generated responses in open-ended contexts. However, further linguistic analyses reveal that models disproportionately rely on negative politeness strategies to create distance even in positive contexts, potentially leading to misinterpretations. While LLMs thus demonstrate an impressive command of politeness strategies, these systematic differences provide important groundwork for making intentional choices about pragmatic behavior in human-AI communication.
@@ 1 Introduction
Speakers do not always say exactly what they mean. For example, we might say a friend’s poem “wasn’t terrible” rather than saying “it was bad” to avoid hurting their feelings Yoon et al. (2020), or just compliment specific elements that we liked without mentioning other elements we didn’t like Goffman (1967); Pinker et al. (2008). These kinds of politeness strategies allow speakers to balance competing goals, conveying accurate information while maintaining positive relationships Hill et al. (1986); Leech (2014). As large language models (LLMs) are increasingly deployed in open-ended interactions across sensitive social domains like healthcare and education, their ability to appropriately use and understand polite language remains an important alignment challenge.
Politeness theory provides a valuable framework for addressing these questions. Seminal work by Brown and Levinson (1987) distinguishes between positive politeness strategies that affirm the listener (compliments, expressions of interest) and negative politeness strategies that minimize imposition (hedging, indirectness). While subsequent work has expanded this framework to encompass broader relational and rapport management concerns (Spencer-Oatey, 2011; Watts, 2003; Locher, 2013), this basic distinction remains crucial: different contexts call for different strategies, and mismatches can lead to communication breakdowns. If an AI system employs negative hedging strategies (“I am somewhat concerned that this approach might not be optimal”) in contexts where human speakers would expect positive, rapport-building strategies (“I love your creativity here, and wonder if we could build on it by…”), users may be left unsure whether the hedging reflects a genuinely negative evaluation or simply a systematic bias in expressions of politeness.
This kind of pragmatic misalignment represents a critical gap in our understanding of LLMs as social agents. While considerable attention has been paid to whether models can recognize politeness or generate polite language in constrained settings, effective social interaction depends not just on understanding politeness norms in the abstract but on actively selecting and applying appropriate strategies from a diverse linguistic repertoire. Human speakers navigate this complexity intuitively, deploying hedging, elaboration, indirect speech acts, and numerous other strategies to balance competing communicative goals in context-sensitive ways. To fully understand LLMs’ grasp of politeness strategies, we need to examine whether they exhibit similar patterns of strategy selection and deployment across different contexts. This requires moving beyond limited-choice evaluations to examine open-ended language generation, where models have access to the full range of linguistic choices. Rather than prescribing how AI systems ought to handle politeness, we seek descriptive evidence about current patterns of behavior.
In this work, we investigate we compare human and LLM politeness strategies in English-language scenarios, making the following contributions:
@@
Our results reveal that while LLMs have acquired important aspects of human-like pragmatic competence in polite language production – enough to be preferred by human evaluators – they also show systematic differences in strategy deployment that raise intriguing questions about the mechanisms underlying their social language capabilities. In particular, we find that models disproportionately rely on negative politeness strategies (minimizing imposition) even in contexts where humans prefer positive politeness strategies (building rapport), suggesting important differences in how these systems navigate social interactions [^1] Data & Code: https://github.com/haoranzhao419/politeness-speech-production.
@@ 2.3 Polite language generation in LLMs
Work on generating polite language in AI systems represents a small but growing research area. Early approaches focused on style transfer, with systems like those developed by Niu and Bansal (2018) demonstrating that neural models could transform neutral text into more polite versions through specific syntactic transformations. Subsequent work explored paraphrasing to increase politeness Fu et al. (2020), politeness-focused style transfer Madaan et al. (2020), and creating polite chatbots Mukherjee et al. (2023). However, these systems typically focused on surface-level transformations rather than strategic deployment of politeness based on contextual factors. As noted in a recent survey Priya et al. (2024), existing approaches to polite language generation have predominantly emphasized isolated features (hedging expressions, please markers, specific lexical choices) rather than examining the full repertoire of politeness strategies and how they’re selected based on communicative context. This leaves a significant gap in our understanding of whether LLMs can approximate the context-sensitivity that characterizes human politeness. Our work addresses this gap by directly comparing politeness strategies in humans and LLMs across varying communicative goals, examining whether models align with human preferences for positive versus negative politeness strategies in different contexts.
@@ 3 Experiment 1: Constrained Settings
Table 1: Spearman correlations between the frequencies of human responses (Yoon et al., 2020) and LLM responses across all goal-rating combinations. Bold values indicate the highest correlation.
To what extent are LLMs sensitive to the goals that give rise to politeness in human speech? To address this question, we first examined whether LLMs could reproduce the patterns of goal-sensitive language use reported by Yoon et al. (2020). Their study provided empirical evidence for a computational model of politeness where speakers strategically balance informational accuracy with social goals. Most notably, they found that when giving negative feedback, humans often deploy indirectness through negation (e.g., “wasn’t terrible” rather than “bad”; see also Gotzner and Scontras, 2024; Lumer and Buschmeier, 2022).
@@
We tested a range of open-source (8B-72B parameters) and closed-source models using two prompting strategies: an “original” strategy that presented scenarios verbatim, and a “persona” variant that systematically varied speaker characteristics (e.g., gender, occupation, background) to better approximate the diversity in the population of human participants Murthy et al. (2025); He-Yueya et al. (2024). For each model and prompting strategy, we sampled 30 responses with temperature $\tau=1.0$ per scenario (see Appendix A for details of prompt design).
@@ 4.1 Methods
We used the scenarios from Yoon et al. (2020), preserving the same performance ratings (0-3 hearts) and communicative goals (informative, social, both, default). We collected 3 open-ended responses per scenario from 156 human participants via Prolific (each responding to 4 distinct scenarios) and three responses from LLMs that performed well in our first experiment: GPT-4o, Claude-3.5-Sonnet, and Llama-3.3-70B-Instruct. We selected these larger models after calibration experiments revealed that smaller models failed to generate coherent responses in the open-ended format (see Appendix B.6). Models were instructed to “keep responses short and concise” to ensure comparable length with human responses.
To assess preferences for these responses, we then conducted a two-alternative forced-choice evaluation with 156 human evaluators, each viewing four different scenarios. Evaluators made five judgments per scenario: (1) comparing human vs. LLM responses, (2-3) comparing goal-congruent vs. goal-incongruent responses for both sources, and (4-5) comparing rating-congruent vs. rating-incongruent responses for both sources. We randomized presentation order and ensured evaluators saw responses from different sources across blocks (see Appendix B for full details).
@@ 5.1 Negation
Category Politeness Strategy Examples
Positive Politeness Gratitude I’m so grateful that …
Be optimistic You can even make them better next time!
Negative Politeness Apologizing I didn’t like it, sorry!
Question or hedge Maybe you can try something different?
Off-Record Be vague It was interesting. (true rating is 0)
Give association clues It was better than those who can’t play.
Bald on-Record Negative lexicon It was terrible!
Factuality I didn’t like the cookies at all.
Table 2: Examples of utterances for different politeness strategies (two per category), with the corresponding sub-strategy highlighted in bold. See the comprehensive list in Table 12.
@@ 5.2 Word usage patterns
To better understand differences between human and LLM responses, we analyzed unigram distributions using Pointwise Mutual Information (PMI) and Jensen-Shannon Divergence (JSD). First, examining unigrams with the highest PMI, we found that human responses more frequently incorporated casual language and expressions (e.g., “awesome," “great"), whereas LLM-generated responses tended toward more formal linguistic choices (e.g., “fabulous," “excellent"). Both groups effectively employed personalization as a politeness strategy, such as directly mentioning the performer’s name (e.g., “Your app is pretty good, Henry!"). Additionally, both humans and LLMs adapted their lexical choices based on context, with minimal overlap in high-PMI words across different goals and ratings.
@@ 5.3 Annotated politeness strategies
To obtain a comprehensive picture of the politeness strategies employed in human and LLM responses, we conducted a detailed annotation using the politeness framework from Brown and Levinson (1987), supplemented by markers from Danescu-Niculescu-Mizil et al. (2013). This framework distinguishes four broad categories of politeness strategies with many subtypes (see Table 2 and Appendix Table 12): positive politeness (e.g. compliments and expressions of interest), negative politeness (e.g. hedging and indirectness), off-record strategies (indirect hints that maintain plausible deniability), and bald-on-record strategies (direct statements without politeness). We used LLMs (GPT-4.1 and Claude-3.7-Sonnet) as annotators, following best practices Tan et al. (2024). Annotations were manually verified and corrected where necessary. We used LLMs (GPT-4.1 and Claude-3.7-Sonnet) as annotators, following best practices Tan et al. (2024). Annotations were manually verified and corrected where necessary. Appendix C.3 provides full details on the annotation process and framework.
Overall strategy distribution.
As shown in Figure 3, both humans and LLMs rely primarily on positive politeness (rapport-building) and negative politeness (minimizing imposition) strategies, with relatively low use of off-record and bald-on-record approaches. However, a key difference emerged in strategy selection patterns: while both humans and LLMs increased their use of positive politeness strategies as ratings increased, LLMs showed systematically higher use of negative politeness strategies even in positive contexts (higher ratings), where humans tended to reduce such strategies. These strategy distributions were significantly different under a permutation test (JSD = 0.023, $p<0.001$). This pattern, where LLMs maintain high levels of hedging and indirectness across contexts, may reflect training objectives that prioritize avoiding potential harm over more human-like affirmation strategies.
@@
Cross-cultural variation.
Given documented cultural differences in politeness norms, we also conducted an exploratory analysis testing whether strategy distributions varied between our US and UK participants. We found a statistically significant difference (JSD = 0.5279, $p<0.001$). US participants deployed a more diverse range of sub-strategies overall than UK participants, especially among positive strategies; however, the effect size was relatively small and our sample was imbalanced (532 US responses vs. 92 UK responses). While this preliminary finding suggests some cross-cultural variation even within English-speaking populations, more balanced sampling would be needed to characterize these differences systematically. Future work should examine how cultural context shapes both human and LLM politeness strategies across a broader range of English varieties and other languages.
6 Discussion
@@
We find that LLMs demonstrate impressive pragmatic competence in capturing politeness phenomena, but also differ in important ways from humans. Most notably, in open-ended production, LLMs continue to rely on negative politeness strategies (hedging, indirectness) even in positive contexts where humans shift toward positive strategies (rapport-building). Interestingly, LLM responses were consistently preferred by human evaluators over crowd-sourced human responses, suggesting that preference judgments may track different features of the utterance than the mere ability to reproduce human-like patterns. This contrast — aligning with human patterns in constrained multiple-choice tasks while diverging in more open-ended tasks — emphasizes that LLMs leverage a richer repertoire of strategies when given the opportunity to express them, such as the familiar “compliment, but criticism” constructions Prochazka et al. (2020) we observed in low-rating scenarios, which were not provided as available options in the constrained tasks.
Recent developments in politeness theory may offer insight into these results. While Brown and Levinson (1987) classically emphasized the role of face threat, subsequent work has reconceptualized politeness as part of a broader class of relational work (Spencer-Oatey, 2011; Locher, 2013). From this perspective, politeness strategies manage the affective quality of relationships through multiple components: face concerns (desires for positive evaluation and acknowledgment of identity), sociality rights (what people expect from each other in interaction, including expectations about appropriate levels of imposition and warmth), and other relational goals.
Our findings suggest that humans and LLMs may functionally assign different weights to these components. Humans emphasize equity rights (avoiding imposition) when delivering criticism but shift toward association rights (building warmth) when giving praise. But LLMs consistently maintain a high weight on imposition across all contexts, potentially aligning with user expectations for AI systems, where maintaining appropriate distance could be more desirable than pursuing interpersonal warmth. Understanding how different training objectives and alignment procedures implicitly shift these relational weights represents an important direction for future work, particularly so developers can make intentional choices about more or less desirable interpersonal dynamics.
Distributional differences in politeness strategies may also have practical implications for human-AI communication. Following Gricean principles of communication as rational social action Grice (1975), listeners expect speakers to modulate their politeness strategies based on the context: you shouldn’t need to take steps to mitigate face threat if face isn’t threatened. So when AI systems violate these expectations, they create an interpretive puzzle: users must determine whether hedged language reflects the literal content being communicated or whether it is an artifact of the system’s persona, risking pragmatic misinterpretations where humans interpret hedged positive feedback as more negative than intended. For example, does “your analysis seems reasonably sound” mean there are specific weaknesses that ought to be addressed, or is this just how the system expresses unqualified approval? Users may adapt their expectations of the system over time (Branigan et al., 2010; Araujo and Bol, 2024; Zhou et al., 2023; Waytz et al., 2014), but may still incur a hidden cognitive cost associated with correcting for these differences. Future work should examine how these differences play out across cultural contexts, where politeness norms already vary substantially (Wierzbicka, 2003; Ide, 1989).
In conclusion, our findings provide an empirical foundation for understanding differences between humans and AI systems in a key domain of pragmatic language use. Critically, our study was descriptive rather than prescriptive. Divergences matter not because artificial simulacra of human-like behavior is desirable, but because it affects human-AI communication in practice, raising questions about how users learn to interpret AI-generated politeness differently than human-generated politeness. Beyond practical applications, these findings also contribute to theoretical debates about the cognitive architecture underlying polite speech, and the computational principles that guide decisions across social contexts. These findings could inform future LLM training approaches to better align with human pragmatic patterns, laying groundwork for making intentional rather than accidental choices about subtle pragmatic patterns in AI systems.
@@
While our work gave a comprehensive picture of comparing the polite language use in humans and LLMs, there are still limitations that could be addressed in future work. First and foremost, we want to point out that our study exclusively focused on English-language politeness strategies. While we acknowledge that politeness strategies vary largely across languages and cultures, cross-linguistic and cross-cultural analyses are beyond the scope of this work and represent important directions for future research.
Regarding the “persona” prompting in Experiment 1, we varied demographic factors (e.g., gender, occupation, etc.) across personas to increase response diversity. While demographic influences on politeness are not the focus of this study, and we observed no clear demographic effects on LLM performance, we acknowledge the potential risks that explicitly using demographic categories in prompts could activate stereotypes in LLM responses. One explanation from a recent study is that specific demographic information can activate LLM stereotypes through “implicit personalization,” where models automatically infer users’ demographic attributes from subtle conversational cues (topics, language patterns, cultural references) and then generate responses based on those stereotypical associations Neplenbroek et al. (2025). To mitigate this risk, future work could explore alternative methods for eliciting response variation, such as varying communicative contexts rather than speaker demographics. Additionally, in our human experiments, we acknowledge that the preferences given by human evaluators were from a third-person perspective, which may underestimate the effects of politeness strategies. For instance, receiving overly hedged criticism (“Your work is a bit lacking”) might feel more threatening or frustrating in a first-person context where it is your own work, while positive rapport-building strategies might be more appreciated by direct recipients. Future research should examine first-person reactions to these politeness patterns, ideally through synchronized interaction scenarios where participants receive feedback from other humans vs. LLMs, which would better capture the relational and emotional dimensions of pragmatic (mis)alignment.
Furthermore, throughout our analyses, we still cannot answer the question of what makes human evaluators prefer the responses they prefer, as all our analyses showed very minimal differences between preferred and non-preferred responses. One guess is that even ratings and goals are made very clear in the provided scenarios, human evaluators still may not pay enough attention to, and optionally omit this information, instead, they tend to pick whichever one in the given pair that sounds nicer. Future research, for example, testing LLMs as evaluators and comparing LLM-as-evaluator preference results with humans, could give us more insight into this question. Also, as our results show that LLMs are still not quite human-like in picking the right politeness strategies in a context-sensitive way, future research on how to develop computational methods and algorithms to make LLMs better at polite language use and as social agents will be necessary.